OpenStudy (anonymous):

how do i set up the equation for this word problem? A solution of 78% alcohol is to be mixed with a solution of 23% alcohol to form 165 L of a 58% solution. How many liters of the 78% solution must be used?

OpenStudy (anonymous):

Let the alcohol with 78% be solution with quantity A. Let the alcohol with 23% be solution with quantity B. 0.78A+0.23B=0.58(165) =95.7 Also the total of A and the total of B must equal 165, in other words A+B=165 So you have simultaneous equations now. Is that enough to get on with?

OpenStudy (anonymous):

0.78 x + 0.23 y = 0.58 * 165 where x = vol. of 78 % solution and y = 23% also x + y = 165

OpenStudy (anonymous):

i gav u the same equations silverstravin....guess u must've gone wrong in calculation. check it again.

OpenStudy (anonymous):

I got x=-0.2948717949y+122.6923076923 ???????????